1. Night Inc. has unlevered cash flows of $1,470,000 each year in perpetuity. The unlevered cost of capital (ru) is 14%. The firm plans to issue $6 million in perpetual debt with a return of 8% (to repurchase stock). The tax rate is 25%. If the value of the levered firm is $11,500,000, use the trade-off theory to find the present value of the financial distress costs.  

Unless stated otherwise, compounding is annual and payments occur at the end of the period.
Answer #1

Value of Unlevered firm =Unlevered Cash flows//Cost of Equity =1,470,000/14% =10,500,000
Value of Debt =6,000,000
Total Levered firm =Value Unlevered +Debt*Tax Rate =10,500,000+6,000,000*25% =12,000,000

Bankruptcy Cost =12,000,000-11,500,000 =500,000